OpenStudy (anonymous):

Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31

OpenStudy (anonymous):

3x-5y=-16 2x+5y=31 So by Adding we get 5x=15.........so.......x=3 so 2(3)+5y=31........so.........5y=25 y=5

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Wlcm :)